In a jib crane, the jib and tie rod are 6 m and 5 m long respectively. The height of the crane post is 3 m. When a weight (W) 317 N is suspended from the crane and the tie rod remains horizontal.


Determine the tension in Tie rod (only magnitude, in N)?


1
Expert's answer
2021-06-04T05:48:03-0400

"\tan\theta = \dfrac{6}{5}"


"\theta = 50.19°"


"\textsf{Tension}= \dfrac{317\ N}{\sin 50.19°}= 412.67\ N"